# SIM2VR: Towards Automated Biomechanical Testing in VR

### v1

@@ 3. The VR Simulation Alignment Problem
As depicted in Figure 1 (right), movement-based interaction in VR can be regarded as a time-continuous closed loop between the two core components of interaction: the user and the VR application. During this interaction, the user obtains multi-sensory information from its environment [^4] Usually, this includes information from the peripersonal and physical space of the user (e.g., proprioceptive or haptic input signals) as well as the virtual space (visual input signals)., then produces a neuromuscular control signal to generate body movements, hence manipulating the input device to complete the given interaction task. The VR application, on the other hand, conveys an interactive, virtual environment that can be manipulated via one or multiple input devices. The state of the input device is actively sensed by the VR system, measuring, e.g., how the VR controller is moved, or whether a physical controller button is pressed. As a final step to close the loop, the VR application provides sensory output to the user, such as a rendered image displayed on an HMD, or haptic output in the form of temporary controller vibrations.
3.1. Aligning Simulated VR Interaction
@@
The VR interaction loop described above not only characterises the movement-based interaction between a real user and a given VR system, but can also be used as a framework for creating model-based simulations of VR interaction. To ensure that the simulated interaction loop acts as a reasonable proxy for real user interaction, i.e., to ensure external validity of the predictions made, the two interaction loops need to be aligned. We coin this problem the VR Simulation Alignment Problem.
@@
As summarised in Table 1 and described in detail below, VR simulation alignment evolves in six dimensions, three of which focus on the alignment of the simulated user and three on the alignment of the simulated VR system.
@@
The physics of the interaction, i.e., the biomechanics of the user’s body and the dynamics of the input and output devices, need to be captured adequately by the simulation. This requires a sophisticated biomechanical model that captures the complexity of the human musculoskeletal system, at least for the body parts relevant for the considered interaction (e.g., a model of the shoulder, arm and hand), as well as simulated input and output devices (e.g., a simulated VR controller and an HMD) that align with the real hardware in terms of static and dynamic properties such as mass or size. Too simple a model of the user’s body (e.g., a linked-segment model of the arm) may produce unrealistic movement trajectories, resulting in severe deviations from how real users interact with a given application.
@@
The simulated VR application should have the same internal states as the target VR application (e.g., it needs to provide the same virtual objects and affordances) and show the same responsive behaviour, given that both applications are provided the same input signals. If application alignment is low, e.g., because a game constraint is missing or is incorrectly re-implemented, applying the same body movements in the simulation as in reality may result in substantially different states of the VR application.
@@
The simulated VR system needs to provide the same output signals as the real VR system, given that both are in the same state. This ensures that the simulated user has the chance to perceive the same information from the application as a real user. In particular, the simulated user needs to be provided the same rendered image as is shown on the the real user’s HMD.
@@
The simulated user should perceive its environment (including its own body, physical input devices and the VR application) in the same way a real user does. This requires appropriate modeling of all sensory modalities relevant for the considered interaction, including vision, proprioception and haptics. Note that perceptual alignment focuses on the observations that users make of their environment, whereas output channel alignment is about the information that the VR application makes available to the user in the first place.
@@
Given an interaction task (e.g., to successfully complete a specific game level as fast as possible), the simulated user should generate the same muscle control signals as a real user, when provided with the same perceptions. In the context of reinforcement learning, this requires defining an appropriate task-specific reward function such that an optimal control policy resembles real user behaviour.
Figure 2. To obtain valid and reliable predictions of interactive user behaviour in VR, a simulated user needs to perceive and control exactly the same VR application as real users. Our SIM2VR system addresses this VR Simulation Alignment Problem by establishing a closed loop between the simulation process and the desired VR application, ensuring that the same input and output signals are generated as with real hardware. The resulting time-continuous integration allows to train and evaluate a biomechanical user simulation directly in a given VR environment.
@@
As the behaviour of the user and the VR application depend on each other in a closed loop manner, errors can easily cascade in time. This results in an increasing divergence between the simulation and the real VR interaction, once an error is introduced. For example, in the Whac-A-Mole game from Section 5, a different grasp (e.g., due to a slight offset in the game dynamics) may result in fewer/more/different targets being hit, which impacts the movements generated by the simulated user at later times, leading to even greater differences in target hits, and so on. The same holds for a missing or incorrectly implemented game constraint: If the simulated user has to hit a target with less power or less often to receive the same game reward, it will learn a totally different control policy, leading to even more divergent application states. Depending on the complexity and the amount of discontinuities in the interaction dynamics, these errors may pollute any variable of interest obtained from the simulation. It is thus essential to ensure alignment in as many dimensions as possible.
4. SIM2VR: Enabling Biomechanical User Simulations in VR
In this section, we propose SIM2VR, a system that allows two simulators, the simulated user and the original VR application, to run in parallel while being synchronised temporally and spatially. The aim of this system is to establish a closed interaction loop between the simulated user and the VR application (see Figure 2) that closely resembles the interaction loop between the real user and the VR application described in Section 3.
In this simulated interaction loop, the simulated user generates muscle control signals based on its sensory perception of the body, the input devices, and the VR application (e.g., visual information of the scene and proprioceptive information of the body posture). These control signals are used to actuate the muscles, driving the movement of the body, and, consequently, the simulated states of the hardware devices. The position and rotation of the VR controllers and HMD are then measured by virtual sensors and sent to the VR application. The VR application processes this input, computes a reward based on the updated application state (e.g., the current game score) and renders the image perceived through the virtual HMD. Finally, the rendered image is provided to the simulated user, along with the reward and optional “stateful” information (such as time remaining in an interaction task), thus closing the interaction loop. [^5] Note that the reward signal and the stateful information are only part of the simulated interaction loop. This is mainly for technical reasons: The reward signal is required for an RL method to learn a control policy, and the stateful information can be used to improve the training.
Crucially, this approach ensures application alignment by giving the simulated user access to exactly the same VR application that real users interact with. Based on this approach, we developed the SIM2VR system, which additionally addresses the input channel and output channel alignment problems to provide more ecologically valid predictions of VR interaction.
@@ 4.1. Implementation Principles & Details
SIM2VR is an extension of the User-in-the-Box (UitB) framework (Ikkala et al., 2022) to Unity applications, which we built based on the following principles outlined in the remainder of this section:
@@
Figure 3. With SIM2VR, simulated users for the first time have access to exactly the same visual output as humans. The top figure depicts how real users perceive a VR environment (in this case, the Whac-A-Mole game from Section 5), while the bottom figure shows the downsampled RGB-D image (120x80 pixels) that the simulated user perceives.
@@
The simulated user can be created within UitB, which requires selecting a biomechanical model implemented in MuJoCo, as well as one or more perception modules. These perception modules define how the simulated user perceives its surroundings, e.g., via visual or proprioceptive signals. To model how users specifically perceive the virtual environment when wearing an HMD, SIM2VR adds a novel vision module denoted as UnityHeadset to UitB. This module enables the simulated user to observe the rendered image sent to the virtual HMD as an RGB-D array (see Figure 3).
In the UitB framework, the interaction environment and the rewards that incentivise completing the given interaction task need to be defined in a so-called task module. To enable using Unity applications as interaction environments, SIM2VR provides a UnityEnv task module on the UitB side, and a SIM2VR Asset on the Unity side. These are connected via ZeroMQ, which creates a communication channel between the two simulators. A detailed overview of the components of SIM2VR and their dependencies is given in Figure 4.
Figure 4. SIM2VR enables efficient data transfer between the simulated user (implemented in UitB) and the VR application (implemented in Unity). This is achieved by establishing a ZeroMQ connection between the generic UnityEnv task module and the SIM2VR Asset, both provided by SIM2VR. The UnityEnv equips the biomechanical model with the VR controller and the HMD. Visual perception of the VR application is received through the UnityHeadset vision module, based on the rendered HMD images sent to the UnityEnv. The SIM2VR Asset also allows application-dependent rewards, such as game scores, to be defined directly in the VR application using the RLEnv Unity script. The rewards can be used by the simulated user to improve its control policy while interacting with the VR application.
@@
By adding the SIM2VR Asset to the desired Unity application, SIM2VR gains direct and genuine access to all Unity variables relevant for the interaction. This has several advantages. First, it permits capturing the output signals, such as the rendered image displayed on the HMD, and directly forwards these signals to the simulated user. [^8] While SIM2VR is currently limited to the transmission of visual output signals, we plan to support other feedback modalities such as auditory and haptic output. It is thus key in ensuring output channel alignment. Second, it can be used to define application-specific rewards, e.g., based on a game score. This is done in the RLEnv script provided by the SIM2VR Asset. Third, it allows to improve the RL training by providing further stateful information (e.g., the time passed in an episode, or the number of completed subtasks) as observation to the simulated user, and to log [^9] To facilitate troubleshooting, we have added Weights & Biases (https://wandb.ai/site) integration, which allows real-time monitoring of the simulated user’s performance. any Unity variables of interest (e.g., the current performance in terms of specific subtasks).
@@
The physics engine of the user simulation and the VR application need to be temporally synchronised. SIM2VR achieves this by including the current as well as the next timestamp of the user simulation as metadata in the data packages sent to Unity. These timestamps are used to halt the VR application updates until the same amount of time has elapsed as in the user simulation. To avoid computational overhead, the Unity scene is rendered exactly once per UitB time step, and only relevant data is transferred between the two simulators. Faster than real time simulations are achieved by modifying the Unity timescale parameter (in our experiments we ran Unity at 5 times real time). The UnityEnv module can also trigger the reset function of the SIM2VR Asset and vice versa. This is essential to ensure that both the simulated user and the VR application are reset together when necessary, for example, to start a new episode during RL training, or after completing a game level.
4.2. Creating the Simulated User
The remaining three alignment problems – physical, perceptual, and cognitive – need to be addressed by improving the quality of the simulated user. While this is not the main focus of SIM2VR, we also provide tools that aim to support VR designers and researchers in this regard.
Reach Envelope Tool
Figure 5. The reach envelope tool provided by SIM2VR visualises the range that the simulated user (black cross, facing in z-axis direction) can theoretically reach with a maximally extended arm (blue lines) along with target positions relevant for the considered VR application (coloured dots). This example shows that in the Whac-A-Mole game (Section 5), all target positions in all three target area placements can be reached by the simulated user, albeit some of the low targets (those close to the boundary of the envelope) require a fully extended arm.
@@ 4.2. Creating the Simulated User
In a 3D plot, the area that can be reached by the simulated user with maximally extended arm is shown as blue lines from a bird’s eye view. A black cross symbolises the position of the simulated user. The figure suggests that the target positions of all three target area placement conditions in the Whac-A-Mole game (visualised by cyan, green and magenta dots for low, mid and high conditions, respectively) are within this reach envelope. Figure title: ”ReachEnvelope Check”
As the biomechanical model typically has fewer degrees-of-freedom (DOFs) than a real human body, it is important to ensure that it is still able to perform all the movements required by the interaction. For a model of the upper extremity used to simulate visuomotor interaction in VR, this mainly involves verifying that all target positions are within the model’s reach envelope. The reach envelope tool provided by SIM2VR allows interactive exploration and visualisation of the body-centred positions that a given simulated user can theoretically reach. This can be used to verify that all target positions that occur in a given VR interaction task can be reached by the simulated user. The tool works as follows. Given a simulated user, its reach envelope is computed by identifying all VR controller positions that can be reached with a maximally extended arm. These positions are then transformed into Unity coordinates and plotted in 3D along with desired target positions (see Figure 5 for an example plot; the tool can also create an animated 360° plot for better spatial perception). Target positions can either be provided manually (in Unity coordinates) or can be directly extracted from a Unity log file. The tool also allows one to distinguish between the envelope of the pure biomechanical model and the envelope when equipped with a VR controller, which can be used to verify correct placement of the controller mesh.
Running this tool for a given simulated user and a given VR interaction task can help improve the physical alignment.
Reward Scaling Tool
Reward functions required to learn biomechanically plausible behaviour typically involve multiple components. For example, effort costs are usually added to the reward to incentivise accomplishing the task with as little muscle effort as possible (Todorov and Jordan, 2002; Croxson et al., 2009). However, balancing these reward terms is a non-trivial task and typically requires some RL experience.
To this end, we provide a reward scaling tool that facilitates reward function design for visuomotor VR interaction tasks. This tool takes the selected reward components (e.g., a one-time bonus reward given at task completion, a dense reward continuously awarded based on the current distance to all active targets, and the effort costs) as input, and then predicts how the contribution of each component to the total reward evolves over time. Predictions are generated by evaluating the reward components for different scenarios such as worst case (i.e., when the simulated user has not learned to make any movements yet), best case (i.e., when the desired target position is already reached at beginning of an episode), and linear or quadratic interpolation between initial and target position. The resulting plots can be used to identify poorly scaled reward components that either have too large of an effect on the simulated user or no effect at all, in which case these components need to be differently weighted.
By helping to identify an appropriate reward function, this tool represents a first step towards better cognitive alignment between simulated and real users, which is certainly the most difficult dimension of the alignment problem to address. A discussion on the current capabilities of simulated users for modelling different VR interaction tasks, and how these might benefit from more sophisticated RL methods recently proposed, can be found in Section 6.

### v2

@@ 3. VR Simulation Alignment
As depicted in Figure 1 (right), movement-based interaction in VR can be regarded as a continuous-time closed loop between the two core components of interaction: the user and the VR application. During this interaction, the user obtains multi-sensory information from their environment and produces a neuromuscular control signal to generate body movements, hence manipulating the input device to complete the given interaction task. The VR application, on the other hand, conveys an interactive, virtual environment that can be manipulated via one or multiple input devices. The state of the input device is actively sensed by the VR system, measuring, e.g., how the VR controller is moved, or whether a physical controller button is pressed. As a final step to close the loop, the VR application provides sensory output to the user, such as a rendered image displayed on an HMD, or haptic output in the form of temporary controller vibrations.
@@
This interaction loop not only characterises the movement-based interaction between a real user and a given VR system, but can also be used as a framework for creating model-based simulations of VR interaction. To ensure that the simulated interaction loop acts as a reasonable proxy for real user interaction, i.e., to ensure external validity of the predictions made, the two interaction loops need to be aligned.
@@
As described in detail below, we categorise VR simulation alignment along six dimensions, three of which focus on the alignment of the simulated user and three on the alignment of the simulated VR system.
@@
The physics of the interaction, i.e., the biomechanics of the user’s body and the dynamics of the input and output devices, need to be captured adequately by the simulation.
@@
The simulated VR application should have the same internal states as the target VR application (e.g., it needs to provide the same virtual objects and affordances) and show the same responsive behaviour, given that both applications are provided the same input signals.
@@
The simulated VR system needs to provide the same output signals (e.g., HMD image) as the real VR system, given that both are in the same state. This ensures that the simulated user has the chance to perceive the same information from the application as a real user.
@@
The simulated user should perceive their environment (including their own body, physical input devices and the VR application) in the same way as a real user would. This requires appropriate modelling of all sensory modalities relevant for the considered interaction, including vision, proprioception and haptics.
@@
Given an interaction task (e.g., to successfully complete a specific game level as fast as possible), the simulated user should have a similar internal representation and select the same control output when given the same sensory input. In the context of reinforcement learning, this requires defining an appropriate task-specific reward function such that an optimal control policy resembles real user behaviour.
As the behaviour of the user and the VR application depend on each other in a closed loop manner, errors can easily cascade in time. This results in an increasing divergence between the simulation and the real VR interaction, once an error is introduced. For example, in the Whac-A-Mole game from Section 5, a different hammer grip (e.g., due to a slight offset in the game dynamics) may result in fewer/more/different targets being hit, which has an impact on the movements generated by the simulated user at later times, leading to even greater differences in target hits, and so on. The same holds for a missing or incorrectly implemented game constraint: if the simulated user has to hit a target with less power or less often to receive the same game reward, they will learn a totally different control policy, leading to even more divergent application states. Depending on the complexity and the number of discontinuities in the interaction dynamics, these errors may pollute any variable of interest obtained from the simulation. It is thus essential to ensure alignment in as many dimensions as possible.
4. SIM2VR: Enabling Biomechanical User Simulations in VR
In this section, we propose sim2vr, a system that allows two simulators, the simulated user and the original VR application, to run in parallel while being synchronised temporally and spatially. The aim of this system is to establish a closed interaction loop between the simulated user and the VR application (see Figure 2) that closely resembles the interaction loop between the real user and the VR application. While aligning the behaviour of the simulated user with that of real users in all dimensions constitutes a highly complex task, sim2vr addresses the system-related alignment problems (input channel, application, and output channel alignment) described in Section 3.
Figure 2. To obtain valid and reliable predictions of interactive user behaviour in VR, a simulated user needs to perceive and control exactly the same VR application as real users. Our sim2vr system addresses this by establishing a closed loop between the simulation process and the desired VR application, ensuring that the same input and output signals are generated as with real hardware. The resulting time-continuous integration allows to train and evaluate a biomechanical user simulation directly in a given VR environment.
@@ 4. SIM2VR: Enabling Biomechanical User Simulations in VR
The core idea is as follows: the simulated user generates muscle control signals based on their sensory perception of the body, the input devices, and the VR application (e.g., visual information of the scene and proprioceptive information of the body posture). These control signals are used to actuate the muscles, driving the movement of the body, and, consequently, the simulated states of the hardware devices. The position and rotation of the VR controllers and HMD are then measured by virtual sensors and sent to the VR application. The VR application processes this input, computes a reward based on the updated application state (e.g., the current game score) and renders the image perceived through the virtual HMD. Finally, the rendered image is provided to the simulated user, along with the reward and optional “stateful” information (such as time remaining in an interaction task), thus closing the interaction loop.
Crucially, this approach ensures application alignment by giving the simulated user access to exactly the same VR application that real users interact with. This is contrary to existing biomechanical frameworks, such as User-in-the-Box (UitB) (Ikkala22) or MyoSuite (caggiano2022myosuite), which are not designed for modelling interaction with real-world VR applications. These frameworks typically run a physics engine that is incapable of capturing the internal states, dynamics, shapes and visuals of a VR application, which fundamentally limits their ability to reliably predict user behaviour. We intentionally refrained from re-implementing the target VR environment in one of these simulation framework for the purpose of this paper, as this would impose considerable implementation and maintenance overheads, and ultimately result in a low-fidelity replica that is inevitably inferior to the VR environment real users are interacting with.
@@ 4.1. Implementation Principles & Details
We have implemented sim2vr as an extension of UitB to Unity [^4] https://unity.com/ applications, based on the following principles (outlined in the remainder of this section):
@@
Figure 3. With sim2vr, simulated users have access to exactly the same visual output as humans. The top figure depicts how real users perceive a VR environment (in this case, the Whac-A-Mole game from Section 5), while the bottom figure shows the downsampled RGB-D image that the simulated user perceives.
@@
The simulated user can be created within UitB, which requires selecting a biomechanical model implemented in MuJoCo, as well as one or more perception modules. These perception modules define how the simulated user perceives their surroundings, e.g., via visual or proprioceptive signals. To model how users perceive the virtual environment when wearing an HMD, sim2vr adds a novel vision module denoted as UnityHeadset to UitB. This module enables the simulated user to observe the Unity scene through the same Camera game object that renders the scene to real users (see Figure 3). The rendered images are presented to the simulated user as 120$\times$80 pixel RGB-D arrays. The module also allows the visual observation to be filtered and enhanced, for example by omitting certain colour channels or stacking prior observations.
In the UitB framework, the interaction environment and the rewards that incentivise completing the given interaction task need to be defined in a so-called task module. To enable using Unity applications as interaction environments, sim2vr provides a UnityEnv task module on the UitB side, and a SIM2VR Asset on the Unity side. These are connected via ZeroMQ, which creates a communication channel between the two simulators. As UnityEnv handles the communication, it first receives the rendered images from the virtual HMD, before relaying them to the UnityHeadset. Later, after the simulated user’s pose has been updated, UnityEnv sends the HMD and controller position and rotation to the SIM2VR Asset. A detailed overview of the components of sim2vr and their dependencies is given in Figure 4.
Figure 4. sim2vr enables efficient data transfer between the simulated user (implemented in UitB) and the VR application (implemented in Unity). This is achieved by establishing a ZeroMQ connection between the generic UnityEnv task module and the SIM2VR Asset, both provided by sim2vr. The UnityEnv equips the biomechanical model with VR controllers and an HMD. Visual perception of the VR application is received through the UnityHeadset vision module, based on the rendered HMD images sent to the UnityEnv. The SIM2VR Asset also allows application-dependent rewards, such as game scores, to be defined directly in the VR application using the RLEnv Unity script. The rewards can be used by the simulated user to improve their control policy while interacting with the VR application.
@@
By adding the SIM2VR Asset to the desired Unity application, sim2vr gains direct and genuine access to all Unity variables relevant for the interaction. This has several advantages. First, it permits capturing the output signals, such as the rendered image displayed on the HMD, and directly forwards these signals to the simulated user. [^7] While sim2vr is currently limited to the transmission of visual output signals, we plan to support other feedback modalities such as auditory and haptic output. It is thus key in ensuring output channel alignment. Second, it can be used to define application-specific rewards, e.g., based on a game score. This is done in the RLEnv class provided by the SIM2VR Asset, which must be inherited and defined separately for each Unity application to ensure compatibility with RL training. Third, it allows the improvement of the RL training by providing further stateful information (e.g., the time passed in an episode, or the number of completed subtasks) as observations to the simulated user, and to log any Unity variables of interest (e.g., the current performance in terms of specific subtasks).
@@
The physics engine of the user simulation and the VR application need to be temporally synchronised. sim2vr achieves this by including the current as well as the next timestamp of the user simulation as metadata in the data packages sent to Unity. These timestamps are used to halt the VR application updates until the same amount of time has elapsed as in the user simulation. To avoid computational overhead, the Unity scene is rendered exactly once per UitB time step, and only relevant data is transferred between the two simulators. Faster than real time simulations are achieved by modifying the Unity timescale parameter (in our experiments we ran Unity at 5$\times$ real time). The UnityEnv module can also trigger the reset function of the SIM2VR Asset and vice versa. This is essential to ensure that both the simulated user and the VR application are reset together when necessary, for example, to start a new episode during RL training, or after completing a game level. The simulators are also spatially synchronised by means of aligning the simulators’ coordinate systems and appropriately transforming coordinates between them.
4.2. Creating User Simulations in Unity: A Step-by-Step Guide
In the following, we first describe how sim2vr can be used to generate user simulations for a given Unity application. Then, following the guide, we demonstrate how sim2vr was used to train a simulated user to play a “Beat Saber” style game implemented in the VR Beats Kit, a third party software freely available on the Unity Asset Store. [^8] https://assetstore.unity.com/packages/templates/systems/vr-beats-kit-168243 A more detailed version of this walkthrough is available at the project website https://github.com/fl0fischer/sim2vr.
4.2.1. Importing the Asset
First, the SIM2VR Asset must be imported into the Unity project. After adding the sim2vr prefab as a game object into the desired scene, the SimulatedUser game object needs to be connected to the VR Controllers and Main Camera provided by the OpenXR plugin. The SIM2VR Asset contains Recorder and Logger game objects, which can be used to record game play and log controller and headset poses, as well as further game data, during training and evaluation. Due to the reliance on game-specific data, the Logger must be adapted to each game separately.
4.2.2. Defining the Reward Function
The majority of reward functions that have been used successfully to date have some common features. Typically, they consist of a task-specific component that incentivises reaching the considered goal(s) (e.g., reaching a target position, maintaining a body posture, or moving at a certain speed) and effort costs. Effort costs usually act as a regularisation, i.e., they constrain the set of possible solutions by favouring policies that achieve the goal with less effort. The task-specific reward can be either sparse (i.e., reward is only given after completing the task) or dense (i.e., rewards are continuously given, depending on how “close” the agent is to completing the task). Dense rewards are preferable from a mathematical point of view, as they provide smooth gradients, guiding the RL agent towards the optimal policy. However, defining the “right” dense rewards for a given interaction task, i.e., rewards that incentivise task completion but are not biased towards a particular strategy, can be quite challenging. We provide a reward scaling tool, which can be used as a starting point for defining such rewards; the tool is described in the supplementary material.
In sim2vr, the task-specific reward needs to be calculated and stored within the CalculateReward method of RLEnv. If the VR application under consideration provides a game score, this score can be used directly to define the task-specific reward. The effort costs can be defined when configuring the UitB user model.
4.2.3. Defining the Reset Functionality
The RLEnv class contains methods that need to be implemented to reset the game to an initial state at the end of each episode. This usually includes destroying game objects created during runtime and resetting variables used to compute the reward. All code related to resetting the reward variables should be defined in the InitialiseReward method. Preparations for the next episode, such as choosing a game level, can be defined in the Reset method, or, if these should be called only once when starting the game, in the InitialiseGame method. The end of an episode is indicated by the UpdateIsFinished method.
4.2.4. Defining the Stateful Information (optional)
Since including an application- and task-dependent time feature as stateful information in the observation can help to train the RL agent, the RLEnv class provides a GetTimeFeature method that allows time information to be appended to the observation of the simulated user. Note that in deterministic games, the inclusion of such a time feature may lead to a control policy that exploits this information instead of relying on (visual) observations.
4.2.5. Defining the Simulated User
After preparing the Unity application for running user simulations, an appropriate user model needs to be created in UitB. This mainly involves the choice of a biomechanical user model (including effort cost) as well as perception modules beyond the default UnityHeadset vision module, both of which can be easily selected within the UitB config file. Other optional parameters in the config file include the position and orientation of the HMD relative to a specific body part of the biomechanical user model, arguments to be passed to the Unity application (e.g., setting game level/difficulty), and RL hyperparameters (e.g., network size, time steps, batch size, etc.).
To ensure that all the target positions required by the VR application can be reached by the biomechanical model, we provide a reach envelope tool. This tool allows interactive visualisation of the body-centred positions that a given simulated user can theoretically reach, along with the target positions that occur in a given VR interaction task; the tool is described in the supplementary material.
@@ 4.2.5. Defining the Simulated User
(a) Game View
(b) Simulated User
Figure 5. a) In the VR Beats game, the player controls two lightsabers and has to cut through boxes that are moving towards them. b) Still frames from evaluations of trained simulated users playing the VR Beats game, about to hit an incoming box. Top: A simulated user trained with a high effort cost weight ($0.05$) has learned to keep their arms down and rely on wrist movements. Bottom: A simulated user trained with a low effort cost weight ($0.001$) moves their arms considerably more, sometimes resulting in implausible postures.
4.2.6. Example: VR Beats Kit
In VR Beats, players wield lightsabers in both hands and attempts to hit all incoming boxes while dodging walls and avoiding mines (see Figure 5a). Since the biomechanical user models available in UitB cannot move their torso, we trained the simulated user to play only the first 10 seconds of the game, which does not include any walls (or mines). During the first 10 seconds, a total of 4 boxes pass the player, and missing a box causes the episode to end. A successful hit is rewarded with a score between 50-200 points, depending on the speed and direction of the hit. Below we describe the definition of each functionality required by sim2vr.
Reward Function
As a task-specific reward, we used the original (sparse) game score (calculated based on the number of correct and incorrect target hits), as it was easy to implement and not prone to unwanted biases. We also used the default neural effort cost implemented in UitB.
Reset Functionality
For this game, it is sufficient to invoke the existing onGameOverEvent to signal that the episode has ended, destroy the existing game objects, and then invoke the onRestart event to restart the game. The reward is initialised in the Reset method. To check if the episode has terminated, we make use of the existing getIsGameRunning method of the VR_BeatManager instance.
Stateful Information
No additional stateful information was provided to the simulated user.
Simulated User
As VR Beats spawns targets for both hands, we implemented a bimanual UitB user model by converting a bimanual version of the MoBL ARMS model [^9] https://simtk.org/frs/?group_id=657 (Saul14) using the MyoConverter (Ikkala20; wang2022myosim). To speed up the training process, we replaced the muscle actuators with torque actuators attached directly to the joints. As perception modules, we used the default UitB proprioception module and the UnityHeadset vision module provided by sim2vr. The former allows the simulated user to infer joint angles, velocities and accelerations, as well as muscle activations and index finger position. The latter is configured to use the red colour and depth channels of the RGB-D image, and to stack the current observation with a delayed (0.2 seconds prior) visual observation; this configuration allows the control policy to distinguish between left and right arm targets, and to infer target movement.
Results
We trained three simulated users with neural effort cost weights $0.001$, $0.01$ and $0.05$, respectively. These were chosen based on the suggestions provided by the reward scaling tool. These cost weights allow us to explore how the behaviour of the simulated user can vary from an “explorative” or “enthusiastic” player to a more “lazy” player. For instance, the simulated user trained with the highest effort cost weight ($0.05$) learned to do only the bare minimum to move their arms and hit the incoming targets. In particular, they hit the targets by relying mostly on wrist movements, suggesting that the game design could benefit from modifying the target box trajectories to encourage larger arm movements. The simulated users trained with lower effort costs learned to move their arms more, with the lowest effort leading to a policy where the arm movements are exaggerated. Figure 5b shows still frames of simulated users trained with either high or low effort cost.